# Transfer Learning Between U.S. Presidential Elections: How Should We Learn From A 2020 Ad Campaign To Inform 2024 Ad Campaigns?

### v1

@@ E.2 Values of County-Level Ad Effect with the OR Estimator
We provide a comprehensive result (i.e., the specific numbers of confidence intervals) for the results presented in Section 6.2. Specifically, Table LABEL:tab:pa.county.ate.voted lists the ad effect estimated by the OR estimator under transportability for each county in Pennsylvania. It also lists the effect direction for the sensitivity analysis after calibration and the smallest $\exp(|\gamma_{1}-\gamma_{0}|)$ that makes the ATE significant. {longtblr}[ caption = County-level ad effect with the OR estimator in Pennsylvania, the smallest $\exp(|\gamma_{1}-\gamma_{0}|)$ that makes the effect significant, and the effect direction. The voting history is coded in the original way as in Aggarwal et al., (2023), i.e., as binary variables of whether one has voted each year from 2000 to 2018., label = tab:pa.county.ate.voted, ] colspec = XX— XX, County $\widehat{\theta}_{\textnormal{OR}}$ (95% CI)×100 Min $\exp\{|\gamma_{1}-\gamma_{0}|\}$ Effect direction
Adams -0.41 (-1.05, 0.22) 1.012072 Negative Only
Allegheny 0.06 (-0.55, 0.67) Neither Direction
Armstrong -0.6 (-1.3, 0.1) 1.006018 Negative Only
Beaver -0.26 (-0.86, 0.34) 1.020201 Negative Only
Bedford -0.77 (-1.55, 0) 1.000000 Negative Only
Berks -0.18 (-0.76, 0.4) 1.022244 Negative Only
Blair -0.52 (-1.19, 0.15) 1.010050 Negative Only
Bradford -0.57 (-1.26, 0.12) 1.006018 Negative Only
Bucks -0.13 (-0.71, 0.46) 1.028396 Negative Only
Butler -0.42 (-1.06, 0.22) 1.014098 Negative Only
Cambria -0.4 (-1.05, 0.24) 1.014098 Negative Only
Cameron -0.54 (-1.24, 0.16) 1.010050 Negative Only
Carbon -0.34 (-0.97, 0.28) 1.016129 Negative Only
Centre -0.06 (-0.62, 0.51) 1.028396 Negative Only
Chester -0.05 (-0.62, 0.52) 1.032518 Negative Only
Clarion -0.57 (-1.28, 0.13) Neither Direction
Clearfield -0.55 (-1.24, 0.14) 1.010050 Negative Only
Clinton -0.45 (-1.1, 0.21) 1.012072 Negative Only
Columbia -0.32 (-0.93, 0.3) 1.016129 Negative Only
Crawford -0.45 (-1.13, 0.22) 1.014098 Negative Only
Cumberland -0.17 (-0.77, 0.43) 1.024290 Negative Only
Dauphin 0.04 (-0.56, 0.63) 1.034585 Negative Only
Delaware 0.02 (-0.59, 0.64) Neither Direction
Elk -0.51 (-1.19, 0.16) 1.010050 Negative Only
Erie -0.19 (-0.78, 0.39) 1.024290 Negative Only
Fayette -0.39 (-1.04, 0.26) 1.014098 Negative Only
Forest -0.61 (-1.34, 0.12) 1.008032 Negative Only
Franklin -0.52 (-1.18, 0.15) Neither Direction
Fulton -0.84 (-1.64, -0.04) 1.000000 Negative Only
Greene -0.49 (-1.15, 0.18) 1.012072 Negative Only
Huntingdon -0.56 (-1.26, 0.15) 1.010050 Negative Only
Indiana -0.39 (-1.04, 0.26) 1.016129 Negative Only
Jefferson -0.63 (-1.36, 0.1) 1.006018 Negative Only
Juniata -0.7 (-1.46, 0.05) 1.004008 Negative Only
Lackawanna -0.09 (-0.71, 0.53) Neither Direction
Lancaster -0.3 (-0.9, 0.29) 1.018163 Negative Only
Lawrence -0.39 (-1.02, 0.25) 1.014098 Negative Only
Lebanon -0.35 (-0.98, 0.27) 1.016129 Negative Only
Lehigh -0.02 (-0.59, 0.56) 1.030455 Negative Only
Luzerne -0.19 (-0.79, 0.42) 1.024290 Negative Only
Lycoming -0.45 (-1.09, 0.2) 1.012072 Negative Only
Mckean -0.54 (-1.22, 0.15) 1.008032 Negative Only
Mercer -0.41 (-1.03, 0.21) 1.012072 Negative Only
Mifflin -0.6 (-1.33, 0.12) Neither Direction
Monroe 0.06 (-0.53, 0.65) Neither Direction
Montgomery 0.03 (-0.58, 0.64) Neither Direction
Montour -0.25 (-0.87, 0.37) 1.020201 Negative Only
Northampton -0.01 (-0.58, 0.56) 1.030455 Negative Only
Northumberland -0.39 (-1.04, 0.26) Neither Direction
Perry -0.5 (-1.22, 0.22) 1.012072 Negative Only
Philadelphia 0.38 (-0.35, 1.11) 1.020201 Positive Only
Pike -0.21 (-0.82, 0.39) 1.020201 Negative Only
Potter -0.67 (-1.43, 0.09) 1.006018 Negative Only
Schuylkill -0.43 (-1.08, 0.21) 1.012072 Negative Only
Snyder -0.58 (-1.27, 0.12) 1.008032 Negative Only
Somerset -0.64 (-1.35, 0.07) 1.004008 Negative Only
Sullivan -0.53 (-1.26, 0.2) 1.012072 Negative Only
Susquehanna -0.37 (-1.08, 0.33) 1.020201 Negative Only
Tioga -0.55 (-1.25, 0.15) 1.008032 Negative Only
Union -0.28 (-0.86, 0.31) 1.018163 Negative Only
Venango -0.51 (-1.19, 0.16) Neither Direction
Warren -0.45 (-1.13, 0.22) Neither Direction
Washington -0.35 (-0.99, 0.28) 1.016129 Negative Only
Wayne -0.41 (-1.08, 0.25) 1.014098 Negative Only
Westmoreland -0.37 (-1, 0.27) 1.016129 Negative Only
Wyoming -0.5 (-1.18, 0.17) 1.010050 Negative Only
York -0.33 (-0.93, 0.28) Neither Direction
E.3 Robustness Checks on County Level Ad Effects
In this section, we provide the estimated county level ad effects from both estimators across three ways of coding the voting history mentioned in Section E.1. Figures E.2 and E.3 show the estimation across these three ways of coding voting history for the OR estimator and the EIF-based estimator, respectively. From each figure, across panels, we find the estimation results are robust to the coding methods, especially for the OR estimator. Comparing Figures E.2 and E.3, we find that the EIF-based estimator generally has wider confidence intervals than the OR estimator for small counties and narrower ones for large counties (county sizes have been listed in Figure E.1).
Figure E.2: Robustness checks of the OR estimator with respect to the voting history coding types. In each panel, the bar represents the ad effect under the transportability assumption. The number is the smallest $\exp(|\gamma_{1}-\gamma_{0}|)$ that makes the ad effect significant, colored by the direction of the corresponding ad effect.
Figure E.3: Estimation results of the EIF-based estimator across the voting history coding types. Panels are organized and scaled in the same way as in Figure E.2.
E.4 Details on Data Pre-Processing for Subgroup Analysis
@@
E.4.1 Percentage of Bachelor’s Degree Or Higher in ZIP Codes
@@
E.4.2 Urbanicty in Census Tracts
@@
E.5 Robustness Checks on Subgroup Analysis in Section 6.3
@@
Figure E.4: Ad effect estimates in subgroups defined by the interaction between gender, urbanicity, and percentage of Bachelor’s degree within the same ZIP-code area using both OR estimator and the EIF-based estimator. The results of the OR estimator are exactly those in Figure 6.3.
@@ Appendix F Replication of Section 6 When Excluding Three States from the Source
Sections F.1 and F.2 provide results for a county-by-county analysis and a subgroup analysis, respectively.
@@ F.1 Ad Effects by Counties
We start with the result with the OR estimator and the original way of coding the voting history in Aggarwal et al., (2023). When $\gamma_{0}=\gamma_{1}=0$, i.e., under transportability, the ad effect is insignificant in all counties of PA. When $\gamma_{1}\gamma_{0}\neq 0$, after calibration, the ad effect is sensitive for a negative effect in 52 counties and insensitive in the other 15 counties. Figure F.1 plots the calibrated county-by-county analysis on the map of PA.
Figures F.2 and F.3 show the analysis result across three types of coding the voting history for the OR estimator and the EIF-based estimator, respectively.
Figure F.1: Sensitivity analysis of the ad effect in every county of PA when the source data only consists of voters in NC and AZ. The color gradient represents the smallest $\exp(|\gamma_{1}-\gamma_{0}|)$ that makes the ad effect significant and negative after calibration. Counties that are insensitive to a significant effect are in gray.
Figure F.2: County-by-county analysis with the OR estimator when the source data consists of NC and AZ across three ways of coding the voting history coding. In each panel, the bar represents the ad effect under the transportability assumption. The number is the smallest $\exp(|\gamma_{1}-\gamma_{0}|)$ that makes the ad effect significant, colored by the direction of the corresponding ad effect.
Figure F.3: County-by-county analysis with the EIF estimator when the source data consists of NC and AZ across three ways of coding the voting history coding. In each panel, the bar represents the ad effect under the transportability assumption. The number is the smallest $\exp(|\gamma_{1}-\gamma_{0}|)$ that makes the ad effect significant, colored by the direction of the corresponding ad effect.

### v2

@@ E.2.1 Numeric Values for Figure 3
We provide a comprehensive result (i.e., the specific numbers of confidence intervals) for the results presented in Section 6.2. Specifically, Table LABEL:tab:pa.county.ate.voted lists the ad effect estimated by the OR estimator under the three cases in parts A ($\Gamma_{0}=\Gamma_{1}=1$) and B ($(\Gamma_{0},\Gamma_{1})=(0.99,1.01)$ and $(\Gamma_{0},\Gamma_{1})=(1.01,0.99)$) of Figure 3 for each county in Pennsylvania. {longtblr}[ caption = County-by-county ad effect with the OR estimator in PA under $(\Gamma_{0},\Gamma_{1})=(1,1)$, $(\Gamma_{0},\Gamma_{1})=(0.99,1.01)$, and $(\Gamma_{0},\Gamma_{1})=(1.01,0.99)$. Each cells lists the TATE with $95\%$ CI in parentheses. , label = tab:pa.county.ate.voted, ] colspec = X— X XX, County $\Gamma_{0}=\Gamma_{1}=1$ $\Gamma_{0}=0.99,\Gamma_{1}=1.01$ $\Gamma_{0}=1.01,\Gamma_{1}=0.99$
Adams -0.41 (-1.05, 0.22) -0.04 (-0.68, 0.59) -0.79 (-1.42, -0.16)
Allegheny 0.06 (-0.55, 0.67) 0.41 (-0.2, 1.03) -0.29 (-0.9, 0.32)
Armstrong -0.6 (-1.3, 0.1) -0.25 (-0.95, 0.45) -0.95 (-1.64, -0.25)
Beaver -0.26 (-0.86, 0.34) 0.09 (-0.51, 0.69) -0.61 (-1.2, -0.01)
Bedford -0.77 (-1.55, 0) -0.44 (-1.21, 0.34) -1.11 (-1.88, -0.34)
Berks -0.18 (-0.76, 0.4) 0.18 (-0.4, 0.76) -0.54 (-1.11, 0.04)
Blair -0.52 (-1.19, 0.15) -0.17 (-0.84, 0.5) -0.87 (-1.53, -0.2)
Bradford -0.57 (-1.26, 0.12) -0.21 (-0.9, 0.48) -0.94 (-1.63, -0.25)
Bucks -0.13 (-0.71, 0.46) 0.22 (-0.36, 0.81) -0.48 (-1.06, 0.11)
Butler -0.42 (-1.06, 0.22) -0.07 (-0.71, 0.57) -0.78 (-1.42, -0.13)
Cambria -0.4 (-1.05, 0.24) -0.05 (-0.7, 0.59) -0.75 (-1.39, -0.11)
Cameron -0.54 (-1.24, 0.16) -0.17 (-0.87, 0.53) -0.91 (-1.6, -0.21)
Carbon -0.34 (-0.97, 0.28) 0.04 (-0.59, 0.67) -0.72 (-1.34, -0.1)
Centre -0.06 (-0.62, 0.51) 0.32 (-0.25, 0.89) -0.43 (-1, 0.13)
Chester -0.05 (-0.62, 0.52) 0.3 (-0.28, 0.87) -0.4 (-0.97, 0.17)
Clarion -0.57 (-1.28, 0.13) -0.23 (-0.94, 0.48) -0.91 (-1.62, -0.21)
Clearfield -0.55 (-1.24, 0.14) -0.2 (-0.89, 0.49) -0.9 (-1.59, -0.21)
Clinton -0.45 (-1.1, 0.21) -0.08 (-0.74, 0.58) -0.81 (-1.47, -0.16)
Columbia -0.32 (-0.93, 0.3) 0.05 (-0.56, 0.67) -0.69 (-1.3, -0.07)
Crawford -0.45 (-1.13, 0.22) -0.1 (-0.78, 0.58) -0.81 (-1.48, -0.13)
Cumberland -0.17 (-0.77, 0.43) 0.19 (-0.41, 0.79) -0.53 (-1.13, 0.06)
Dauphin 0.04 (-0.56, 0.63) 0.41 (-0.19, 1) -0.33 (-0.93, 0.26)
Delaware 0.02 (-0.59, 0.64) 0.38 (-0.24, 1) -0.33 (-0.95, 0.29)
Elk -0.51 (-1.19, 0.16) -0.15 (-0.83, 0.53) -0.87 (-1.55, -0.2)
Erie -0.19 (-0.78, 0.39) 0.16 (-0.43, 0.74) -0.54 (-1.12, 0.04)
Fayette -0.39 (-1.04, 0.26) -0.02 (-0.67, 0.63) -0.76 (-1.4, -0.11)
Forest -0.61 (-1.34, 0.12) -0.26 (-0.99, 0.47) -0.97 (-1.7, -0.24)
Franklin -0.52 (-1.18, 0.15) -0.16 (-0.83, 0.51) -0.87 (-1.54, -0.21)
Fulton -0.84 (-1.64, -0.04) -0.5 (-1.3, 0.31) -1.19 (-1.98, -0.39)
Greene -0.49 (-1.15, 0.18) -0.13 (-0.8, 0.54) -0.84 (-1.51, -0.17)
Huntingdon -0.56 (-1.26, 0.15) -0.21 (-0.91, 0.5) -0.91 (-1.61, -0.2)
Indiana -0.39 (-1.04, 0.26) -0.02 (-0.68, 0.63) -0.75 (-1.39, -0.1)
Jefferson -0.63 (-1.36, 0.1) -0.27 (-1, 0.46) -0.99 (-1.71, -0.26)
Juniata -0.7 (-1.46, 0.05) -0.38 (-1.13, 0.38) -1.03 (-1.78, -0.28)
Lackawanna -0.09 (-0.71, 0.53) 0.27 (-0.35, 0.89) -0.44 (-1.06, 0.17)
Lancaster -0.3 (-0.9, 0.29) 0.04 (-0.55, 0.64) -0.65 (-1.24, -0.06)
Lawrence -0.39 (-1.02, 0.25) -0.03 (-0.66, 0.61) -0.75 (-1.39, -0.12)
Lebanon -0.35 (-0.98, 0.27) 0.01 (-0.61, 0.64) -0.72 (-1.34, -0.1)
Lehigh -0.02 (-0.59, 0.56) 0.36 (-0.21, 0.94) -0.39 (-0.97, 0.18)
Luzerne -0.19 (-0.79, 0.42) 0.17 (-0.43, 0.78) -0.55 (-1.15, 0.05)
Lycoming -0.45 (-1.09, 0.2) -0.1 (-0.74, 0.55) -0.8 (-1.44, -0.15)
McKean -0.54 (-1.22, 0.15) -0.16 (-0.85, 0.52) -0.91 (-1.59, -0.23)
Mercer -0.41 (-1.03, 0.21) -0.05 (-0.67, 0.57) -0.77 (-1.38, -0.15)
Mifflin -0.6 (-1.33, 0.12) -0.26 (-0.98, 0.47) -0.95 (-1.67, -0.23)
Monroe 0.06 (-0.53, 0.65) 0.46 (-0.13, 1.05) -0.34 (-0.92, 0.25)
Montgomery 0.03 (-0.58, 0.64) 0.38 (-0.22, 0.99) -0.32 (-0.93, 0.28)
Montour -0.25 (-0.87, 0.37) 0.14 (-0.48, 0.76) -0.63 (-1.25, -0.01)
Northampton -0.01 (-0.58, 0.56) 0.37 (-0.2, 0.94) -0.4 (-0.97, 0.17)
Northumberland -0.39 (-1.04, 0.26) -0.02 (-0.67, 0.63) -0.76 (-1.41, -0.11)
Perry -0.5 (-1.22, 0.22) -0.15 (-0.87, 0.58) -0.86 (-1.58, -0.14)
Philadelphia 0.38 (-0.35, 1.11) 0.75 (0.02, 1.48) 0.01 (-0.71, 0.74)
Pike -0.21 (-0.82, 0.39) 0.19 (-0.41, 0.79) -0.62 (-1.22, -0.01)
Potter -0.67 (-1.43, 0.09) -0.32 (-1.08, 0.44) -1.03 (-1.79, -0.27)
Schuylkill -0.43 (-1.08, 0.21) -0.07 (-0.72, 0.58) -0.79 (-1.44, -0.15)
Snyder -0.58 (-1.27, 0.12) -0.23 (-0.93, 0.47) -0.92 (-1.62, -0.23)
Somerset -0.64 (-1.35, 0.07) -0.31 (-1.02, 0.41) -0.97 (-1.68, -0.26)
Sullivan -0.53 (-1.26, 0.2) -0.17 (-0.9, 0.56) -0.88 (-1.61, -0.16)
Susquehanna -0.37 (-1.08, 0.33) -0.01 (-0.72, 0.7) -0.73 (-1.44, -0.03)
Tioga -0.55 (-1.25, 0.15) -0.19 (-0.89, 0.51) -0.91 (-1.61, -0.21)
Union -0.28 (-0.86, 0.31) 0.1 (-0.48, 0.69) -0.65 (-1.23, -0.07)
Venango -0.51 (-1.19, 0.16) -0.16 (-0.83, 0.51) -0.87 (-1.54, -0.2)
Warren -0.45 (-1.13, 0.22) -0.09 (-0.76, 0.58) -0.82 (-1.49, -0.15)
Washington -0.35 (-0.99, 0.28) 0.01 (-0.62, 0.65) -0.72 (-1.35, -0.09)
Wayne -0.41 (-1.08, 0.25) -0.05 (-0.71, 0.62) -0.78 (-1.44, -0.12)
Westmoreland -0.37 (-1, 0.27) -0.02 (-0.65, 0.62) -0.72 (-1.35, -0.09)
Wyoming -0.5 (-1.18, 0.17) -0.15 (-0.83, 0.52) -0.85 (-1.53, -0.18)
York -0.33 (-0.93, 0.28) 0.03 (-0.58, 0.63) -0.68 (-1.29, -0.08)
E.2.2 Additional Example Sensitivity Parameters
To supplement part B of Figure 3, Figure E.2 plot the conclusions under additional choices of $(\Gamma_{0},\Gamma_{1})$s.
Figure E.2: Conclusions of ad effects with given sensitivity parameters.
E.3 Comparison Between Estimators
In this section, we compare results from the the two estimators: the OR estimator as presented in Section 6.2 of the main text and the EIF-based estimator. Figure E.3 compares the results under transportability across the three ways of coding the voting history (see Section E.1 for the coding of voting history). In each panel, the x- and y-axes represent the results from the OR estimator and the EIF estimator; we find the points lie around the $y=x$ line which indicates that the point estimates are close. The CIs also have comparable lengths. Numeric values of the estimates (CIs) are given in Figures E.4 and E.6 for the OR and EIF-based estimators, respectively.
Figure E.3: Comparison between the OR and EIF-based estimators for estimating ad effects for every PA county under transportability. Each panel represents one way of coding the voting history; the left panel corresponds to the way presented in the main text. In each panel, x- and y- axes represent results from the OR and the EIF-based estimator, respectively. The points represent point estimates and gray bars represent 95% CIs. The dashed line represents $x=y$.
We zoom into the first way of coding the voting history, which corresponds to Aggarwal et al., (2023) and the main text; results for the other two ways are presented in Sections E.4 and E.5 for the OR and EIF-based estimators, respectively. When $\Gamma_{0}=\Gamma_{1}=1$, the EIF-based estimator yields no significant results while the OR estimator yields a significant and negative effect in Fulton. The conservative result from the EIF-based estimator is due to the small sample size of voters in the Fulton county ($n_{t}=4746$): the small $n_{t}$ leads to finite-sample violations to Assumption 2.2 and difficulty in estimating the density ratio $w({\bf V})$, which in turn yields higher variance estimates. We note that this phenomenon only happen for a few small counties, and in general, the widths of CIs for one estimator do not uniformly dominate the other.
When $\Gamma_{0}\neq 1$ or $\Gamma_{1}\neq 1$, we apply the calibration procedure for each estimator following the same procedure as in Section 6.2. Within the calibrated region, the analysis based on the EIF-based estimator produced nine more counties sensitive for a positive effect and two fewer counties sensitive for a negative effect than those in Section 6; see Figures E.5 and E.7.
E.4 Robustness Checks on County Level Ad Effects for the OR Estimator
In this section, we check robustness of the OR estimator across three ways of coding the voting history mentioned in Section E.1. From Figure E.4, we find the 95% CIs under transportability are similar across ways of coding the voting history with slight differences. The second way of coding the voting history gives no significance while the third way yields significance for Fulton, Potter and Bedford counties which all have a negative effect.
Figure E.5 provides the conclusions under transportability on the top panels and the changes in conclusions within the calibrated sensitivity analysis on the bottom panels. We note that the top left and bottom left panels are parts A and D of Figure 3 in the main text. For the calibrated sensitivity analysis, the results of 51 counties are the same across the three ways of coding voting history. It’s notable that the Philadelphia county continues to be only county sensitive for a positive effect. For the other 16 counties, discrepancy occurs mainly due to the differences in significance under transportability. It’s notable that the Monroe county can be insensitive, sensitive for a negative effect, sensitive for a positive effect depending on the way of coding the voting history.
Figure E.4: Robustness checks of the OR estimator with respect to the voting history coding types. Each panel plots the 95% CI under transportability on the x-axis with color indicating significance.
Figure E.5: Robustness checks of the OR estimator with respect to the voting history coding types. The top row represents results under transportability and the bottom row represents the change of conclusions after the calibrated sensitivity analysis.
E.5 Robustness Checks on County Level Ad Effects for the EIF-Based Estimator
In this section, we check robustness of the EIF-Based estimator across three ways of coding the voting history mentioned in Section E.1. From Figure E.6, we find the 95% CIs under transportability are similar across ways of coding the voting history. The ad effects are insignificant except for one case: the ad effect is significant and negative in Juniata county in the second way of coding the voting history.
Figure E.7 provides the conclusions under transportability on the top panels and the changes in conclusions within the calibrated sensitivity analysis on the bottom panels. For the calibrated sensitivity analysis, the results are consistent in 45 counties across ways of coding voting history. In the 21 counties with inconsistent results, we find the highest discrepancy in the Centre county and the Chester county. The Centre county is sensitive to a significant effect of either sign, to a significant and negative effect, and insensitive in the three ways of coding. The Chester county is sensitive to a significant and positive effect, to a significant and negative effect, and insensitive in the three cases.
Figure E.6: Robustness checks of the EIF estimator with respect to the voting history coding types. Each panel plots the 95% CI under transportability on the x-axis with color indicating significance.
Figure E.7: Robustness checks of the EIF estimator with respect to the voting history coding types. The top row represents results under transportability and the bottom row represents the change of conclusions after the calibrated sensitivity analysis.
E.6 Details on Data Pre-Processing for Subgroup Analysis
@@
E.6.1 Percentage of Bachelor’s Degree Or Higher in ZIP Codes
@@
E.6.2 Urbanicty in Census Tracts
@@
E.7 Robustness Checks on Subgroup Analysis in Section 6.3
@@
Figure E.8: Ad effect estimates in subgroups defined by the interaction between gender, urbanicity, and percentage of Bachelor’s degree within the same ZIP-code area using both OR estimator and the EIF-based estimator. The results of the OR estimator are exactly those in Figure 5.
@@ Appendix F Replication of Section 6 When Excluding Three States from the Source
Sections F.1 and F.2 provide results for a county-by-county analysis and a subgroup analysis, respectively, which mirror Sections 6.2 and 6.3 of the main text.
@@ F.1 Ad Effects by Counties
Figures F.1 and F.2 plot the results from the OR estimator and the EIF-based estimator, respectively. When $\Gamma_{0}=\Gamma_{1}=1$, i.e., under transportability, the ad effect is insignificant in all counties of PA for both estimators. When $\Gamma_{0}\neq 1$ or $\Gamma_{1}\neq 1$, after calibration, the ad effect is sensitive for a negative effect in 42 counties from the OR estimator and 10 counties from the EIF estimator. Results in the other counties are insensitive. We note that the result while restricting the source with data from NC, AZ alone give more conservative result than Section 6 due to the smaller sample size of the source population.
Figure F.1: County-by-county analysis with the OR estimator when the source data only consists of voters in NC and AZ. The left panel represents the insignificance of the result under transportability. In the right panel, the red downward arrow represents counties sensitive to a signifcant and negative effect.
Figure F.2: County-by-county analysis with the EIF estimator when the source data only consists of voters in NC and AZ. The left panel represents the insignificance of the result under transportability. In the right panel, the red downward arrow represents counties sensitive to a signifcant and negative effect.